• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jmaxwell

    Log entry time 19:03:49 on March 15, 2012

    Entry number 366585

    keyword=[Target] Buffer Dewar Valve Work Finished

    The target group made significant improvements to the target buffer dewar valve. The optical encoder was replaced with a potentiometer, which should be less sensitive to the radiation which was causing the read errors.

    In addition, the transfer line it self was pulled and the valve stem removed. Additional O-rings were added to the stem to reduce the chance of a leak. The stem and transfer line were returned and the cryo should be back to normal. The TO will have to put the PID loop on the run valve back off manual as soon as the liquid level on the nose is reading again, as it is now over level.